OpenStudy (thatonegirl_):

Approximately 4% of people are colorblind. Consider a class of 460. What is the probability that 15 or fewer students are colorblind?

OpenStudy (anonymous):

do you want a fraction?

OpenStudy (perl):

This problem can be solved using a binomial distribution.

OpenStudy (anonymous):

oh i obviously dont belong here then ....;-; carry on young chap

OpenStudy (perl):

$$ \Large \Pr(X \le k) = \sum_{i=0}^{k} {n\choose i}p^i(1-p)^{n-i} \\ \therefore \\ \Large \Pr(X \le 15) = \sum_{i=0}^{15} {460\choose i}0.4^i(1-0.4)^{460-i} $$

OpenStudy (anonymous):

wow you guys are smart o,o this is like next gen math 4 me

OpenStudy (thatonegirl_):

What o.o @perl

OpenStudy (thatonegirl_):

I know it's a binomial distribution and first i find the mean which wold be 18.4 and the standard deviation is 4.2, right?

OpenStudy (perl):

you would do that if you are going to approximate the binomial with a normal standard curve. but the problem did not say to do this

OpenStudy (perl):

want to do it that way?

OpenStudy (thatonegirl_):

Well when n is sufficiently large enough can't you consider it to be normal?

OpenStudy (thatonegirl_):

That's the only thing I've learned Dx

OpenStudy (perl):

yes you can, it just didnt say to use normal approximation explicitly . we could solve this exactly using a calculator

OpenStudy (thatonegirl_):

Oh so you don't need to find the z value?

OpenStudy (perl):

normal curve binomial approximation was useful before the age of calculators, since solving 450 choose 15 is difficult by hand

OpenStudy (perl):

well lets do it both ways, and we can compare the results

OpenStudy (thatonegirl_):

kk

OpenStudy (perl):

there is a typo above in the formula it should be $$ \Large \Pr(X \le 15) = \sum_{i=0}^{15} {460\choose i}\color{red}{0.04}^i(1-\color{red}{0.04})^{460-i} $$

OpenStudy (thatonegirl_):

Oh okay.. Is that a binomial setting just with that first thing?

OpenStudy (perl):

right, the exact answer is `0.251193749212680812548679303635273122372737486042499189442302...` now lets compare this to the normal approximation

OpenStudy (perl):

mean = n*p st. dev = sqrt( n * p * (1-p)) mean = 460 * .04 = 18.4 st.dev = sqrt( 460 * .04 * .96) = 4.202856 First find the z score, for X = 15, since we want less than it. also make a continuity correction because we want P ( X <= 15) Z = ( 15.5 - 18.4) / ( 4.202856)

OpenStudy (thatonegirl_):

.209?

OpenStudy (perl):

Z = -.6900

OpenStudy (thatonegirl_):

ohh i thought u mean on the chart.

OpenStudy (thatonegirl_):

.2451?

OpenStudy (perl):

oh ok lets check the chart . but do you agree with with that z score? P( X <= 15) = P( Z <= -.6900) = 0.245094

OpenStudy (perl):

yes i got the same

OpenStudy (thatonegirl_):

yeah

OpenStudy (perl):

and thats pretty close to the exact answer

OpenStudy (thatonegirl_):

so the answer is 24.51%?

OpenStudy (thatonegirl_):

oh cool

OpenStudy (perl):

yes, using the normal approximation

OpenStudy (thatonegirl_):

How would i do probability of more than 12? with the same data

OpenStudy (thatonegirl_):

12 or more*

OpenStudy (thatonegirl_):

I got a z-score of -1.52 but you would use 1.52 because it's greater than, right?

OpenStudy (perl):

right

OpenStudy (thatonegirl_):

I end up getting 93.57% which is very off

OpenStudy (perl):

so if you are using a table, once you find the area to the left of the z score you can subtract from 1 to get the area to the right of that z score

OpenStudy (thatonegirl_):

So if i did -1.52, just find that value and subtract it from 1?

OpenStudy (perl):

i think we need to make another continuity correction

OpenStudy (perl):